OpenStudy (mortonsalt):

Hello! I was wondering if someone can help me with this question? (Will be posted below.)

OpenStudy (mortonsalt):

Find the limit as it approaches infinity when \[\frac{4x-1}{x} < f(x) < \frac{4x^2+3x}{x^2}\] for all x>5.

OpenStudy (anonymous):

you need to find out what x is?

OpenStudy (anonymous):

or you need the answer?

OpenStudy (mortonsalt):

Im just trying to figure out where to begin. :) Thank you!

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

He will help you a lot then

OpenStudy (mortonsalt):

Is it 4?

OpenStudy (anonymous):

no

OpenStudy (welshfella):

sorry cant help with this one

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

is this a multiple choice question?

OpenStudy (mortonsalt):

Nope!

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

10x^2

OpenStudy (mortonsalt):

Alright, thanks!

OpenStudy (anonymous):

:D

zepdrix (zepdrix):

Take limit of left side,\[\large\rm \lim\limits_{x\to\infty}\frac{4x-1}{x}=a\]It will approach some value. Take limit of the other side,\[\large\rm \lim\limits_{x\to\infty}\frac{4x^2+3x}{x^2}=a\]You'll find that they approach the same value! So by Squeeze Theorem: \(\large\rm f(x)\to a\) as \(\large\rm x\to\infty\). It's a little disturbing that you have strict inequalities.. those shouldn't be strict. Maybe a typo? Anyway, yes 4 sounds right! :)